OpenStudy (anonymous):

solve 0.4x+4<=0.8x-9

OpenStudy (angela210793):

0.4x+4<=0.8x-9 0.4x-0.8x<=-9-4 -0.4x<=-13(*-1) 0.4x>=13 x>=13/0.4 x>=32.5